# PixelVLA: Advancing Pixel-level Understanding in Vision-Language-Action Model

### v1

@@ 4.2 Visuomotor Tuning Data Generation
Gripper-aware Region Proposal Stage. Given a sequence of observations $\{\mathbf{x}^{1}_{\eta},\mathbf{x}^{2}_{\eta},\dots,\mathbf{x}^{N_{\eta}}_{\eta}\}$ from the $\eta$-th episode, the first gripper-close state in the episode as $G_{\eta}\in\{1,2,\dots,N_{\eta}\}$ and the corresponding observation as $\mathbf{x}^{G_{\eta}}_{\eta}\in\{\mathbf{x}^{1}_{\eta},\mathbf{x}^{2}_{\eta},\dots,\mathbf{x}^{N_{\eta}}_{\eta}\}$. Here, $N_{\eta}$ represents the length of the $\eta$-th episode. Sequentially, we can select a series of gripper-close state observations $\{\mathbf{x}^{G_{1}}_{1},\mathbf{x}^{G_{2}}_{2},\dots,\mathbf{x}^{G_{N_{e}}}_{N_{e}}\}$ from the whole dataset, where $N_{e}$ denotes the number of total episodes in the dataset. Furthermore, we assume $\{\mathbf{x}^{G_{1}}_{1},\mathbf{x}^{G_{2}}_{2},\dots,\mathbf{x}^{G_{N_{e}}}_{N_{e}}\}$ as a discrete video and apply SAM 2 (Ravi et al. (2024)) to detect the gripper to generate $N_{e}$ region proposals $\{\mathbf{R}_{1},\mathbf{R}_{2},\dots,\mathbf{R}_{N_{e}}\}$, where $\mathbf{R}_{\eta}\in\mathbb{R}^{4}$ is the region proposal for the $\eta$-th episode. In this way, the region proposals can be leveraged to accurately capture object regions from cluttered and low-quality robot observations.
@@
Multimodal Object Segmentation Stage. Given a manipulation instruction such as “Put the Eggplant in Yellow Basket”, we employ a LLM to reason over the instruction and extract the textual description of the target object to be manipulated, e.g., “Eggplant”. For the $\eta$-th episode, we then provide the target object text along with the region proposal $\{\mathbf{R}_{\eta}\}$ into an open-vocabulary object detector Grounding DINO (Liu et al. (2024)) and SAM (Kirillov et al. (2023)). These models detect all relevant object instances, generate their mask annotations, and associate them with the corresponding language expressions from the target object text. We then filter the predictions based on their confidence scores, retaining only the mask annotations within the bounding box that has the highest box-confidence. Sequentially, we derive visual prompts from the object masks by randomly sampling points within the mask, generating random lines inside the object region, and extracting external bounding boxes through mask contour detection.
Finally, we apply the proposed two-stage automated annotation pipeline to the publicly available Fractal dataset (Brohan et al. (2022)) and Bridge v2 dataset (Walke et al. (2023)). From over 280K transitions, we filter out approximately 19.2% failed samples caused by incorrect region proposals or object detections. The resulting dataset, Pixel-160K, contains 160K robot manipulation episodes and 6.5M image–text–action triplets with visual prompts and mask annotations.
@@ 4.3 Visuomotor Instruction Tuning Procedure
To advance fine-grained pixel-level understanding in VLAs, we propose a novel visuomotor instruction tuning procedure, consisting of a continuous action training stage and a pixel-level understanding enhancement stage. Concretely, the first continuous action training stage enables the VLA model to acquire robust continuous action representations from a large mixture of image–text–action datasets. In the second stage, pixel-level understanding is explicitly enhanced by adapting the pretrained model on Pixel-160K dataset through LoRA adaptation (Hu et al. (2022)). The following sections elaborate on the key designs of this two-stage training strategy.
Continuous Action Training Stage. Before training, we initialize the vision encoder, the MLP projector, and the LLM backbone in PixelVLA with the pretrained weights of VLAs (Kim et al. (2024); Black et al. (2024)), which has been trained on the large-scale mixture dataset OXE (O’Neill et al. (2024)). In addition, during this stage, the visual prompting encoder and the multiscale pixel-aware encoder of PixelVLA are removed, while all other modules except the continuous action decoder are frozen to preserve the general manipulation knowledge learned in the pretrained VLAs. To directly map the final hidden states of the last layer of LLM to continuous action values, we follow (Zhao et al. (2023); Kim et al. (2025)) to implement L1 regression to align predicted actions generated by the proposed continuous action decoder with the ground-truth actions. Similar to OpenVLA, we uniformly discretize each continuous action representation into 256 bins and normalize each action dimension to the range [1,+1][-1,+1] for fine-grained action prediction. Furthermore, during this stage, we train PixelVLA on a mixture of Fractal dataset and Bridge v2 dataset.
Pixel-level Understanding Enhancement Stage. Originally, most existing visuomotor instruction tuning methods (Li et al. (2025a); Kim et al. (2025); Yang et al. (2025)) focus on image-level understanding. In contrast, at this stage, to enhance pixel-level understanding of PixelVLA, we employ LoRA adaptation to efficiently fine-tune PixelVLA’s LLM backbone on Pixel-160K dataset, while jointly training the visual prompting encoder along with the multiscale pixel-aware encoder. Meanwhile, the continuous action decoder is optimized while the remaining PixelVLA modules remain frozen. Furthermore, we adopt the same L1 regression loss and continuous action representation strategy as those employed in the continuous action training stage. At each training step, given a mini-batch $\{\mathbf{x}^{i},\mathbf{p}^{i},\mathbf{a}^{i},\mathbf{L}^{i},\mathbf{V}^{i}\}^{B}_{i=1}$ sampled from the Pixel-160K dataset, the forward process at a single timestep of this stage can then be formulated as follows:
5 Experiments
Table 1: SimplerEnv (Li et al. (2024b)) simulation valuation results in terms of the average success rate for the Google Robot setup. VM denotes Visual Matching and VA is Variant Aggregation. ◼ and ◼ denote tuning-based methods applied to the pretrained weights of OpenVLA and $\pi_{0}$, respectively.
Methods
Pick Coke Can Move Near Open/Close Drawer Average
@@ 5.1 Experimental Setup
Evaluation tasks. We conduct all experiments on three simulation benchmarks, i.e., SimplerEnv-Google Robot (Li et al. (2024b)), SimplerEnv-WidowX (Li et al. (2024b)) and LIBERO (Liu et al. (2023a)). SimplerEnv (Li et al. (2024b)) is an open-source simulation suite that facilitates reproducible and scalable evaluation of robot manipulation policies by explicitly addressing visual and dynamic gaps between simulation and real hardware. In light of this, we conduct zero-shot object manipulation comparisons on SimplerEnv. In addition, following OpenVLA (Kim et al. (2024)), we evaluate performance of new robot adaptation across four task suites within LIBERO (Liu et al. (2023a)), i.e., LIBERO-Spatial, LIBERO-Object, LIBERO-Goal and LIBERO-Long.
Implementation Details. To evaluate the effectiveness of PixelVLA, we apply its architecture and the proposed visuomotor instruction-tuning procedure to two widely-used VLAs, OpenVLA (Kim et al. (2024)) and $\pi_{0}$ (Black et al. (2024)). Regarding the training data, PixelVLA is trained in two stages: the first stage utilizes real-robot demonstrations from the Fractal dataset (Brohan et al. (2022)) and Bridge v2 dataset (Walke et al. (2023)), while the second stage employs 160K real-robot demonstrations from the proposed Pixel-160K dataset. For input robot observations across all datasets, PixelVLA is conditioned solely on a single third-person camera view and processes images at a resolution of 224×224 pixels. In all training stages, we set action chunk size to 8 for the continuous action decoder, i.e., the predicted action $\mathbf{a^{t}}\in\mathbb{R}^{8\times 7}$. The first training stage involves training PixelVLA for 100k steps with a batch size of 32 and a learning rate of $5\times 10^{-4}$. Notably, in light of the effectiveness action expert in $\pi_{0}$, we omit the first training stage when adapting PixelVLA on $\pi_{0}$. During the second training stage, we fine-tune the LLM backbone of PixelVLA using LoRA adaptation with a rank $r=32$. This stage is trained for 200k steps with a batch size of 32 and a learning rate of $1\times 10^{-3}$. In addition, to adapt PixelVLA to the LIBERO benchmark (Liu et al. (2023a)), we fine-tune the pre-trained model for 150K steps on each task suite using LoRA
adaptation with rank $r=32$, a batch size of 32, and a learning rate of $5\times 10^{-4}$. In addition to the two baseline VLAs, OpenVLA (Kim et al. (2024)) and $\pi_{0}$ (Black et al. (2024)), we compare the performance of PixelVLA against other state-of-the-art VLAs, such as RT-1 (Brohan et al. (2022)), HPT (Wang et al. (2024)), Octo (Team et al. (2024)), TraceVLA (Zheng et al. (2024)), RoboVLMs (Liu et al. (2025)), Dita (Hou et al. (2025)) and SpatialVLA (Qu et al. (2025)).
@@
Methods
Put Spoon Put Carrot Stack Blocks Put Eggplant Average
@@ Methods
5.2 Zero-shot Object Manipulation Comparisons
This subsection evaluates the zero-shot manipulation performance of our model against baseline VLAs across multiple task categories and robot platforms. As shown in Tab. 1, in the Google Robot setup, PixelVLA achieves a notable average Visual Matching (VM) score of 61.4 and Variant Aggregation (VA) score of 50.1, substantially exceeding the OpenVLA baseline by margins of 28.7 and 10.1 in VM and VA, respectively. These results indicate a strong capability in both pixel-level understanding and adaptation to textual and visual prompts in out-of-domain adaptation. Notably, as shown in Fig. 4, PixelVLA outperforms TraceVLA and OpenVLA across various environmental variations, highlighting the effectiveness of the proposed visuomotor instruction tuning procedure in addressing out-of-domain generalization.
Table 3: Quantitative ablation studies on Variant Aggregation for the Google Robot setup, evaluated in the SimplerEnv simulation environment (Li et al. (2024b)).
Methods
Pick Move Open/Close
Average
 Coke Can  Near  Drawer
Baseline 54.5 47.7 17.7 40.0
       +FT 51.9 42.3 16.8 37.0
       +FT+CAT 61.3 52.3 17.7 43.8
       +FT+PUE 71.1 54.7 21.3 48.0
PixelVLA 72.7 57.7 20.0 50.1
5.3 New Robot Setups Adaptation Comparisons
To evaluate the adaptability of PixelVLA to novel robotic setups and task configurations, we employ the proposed automated annotation pipeline to process the LIBERO benchmark training data (Liu et al. (2023a)), yielding the LIBERO-Pixel dataset. Subsequently, as summarized in Table III, PixelVLA achieves state-of-the-art performance with an average success rate of 86.7 across all tasks, significantly surpassing strong baselines. In addition, PixelVLA surpasses the baselines OpenVLA and TraceVLA by $10.2\%\sim 8.6\%$ in terms of average success rate, respectively. These superior results demonstrate the enhanced adaptability of PixelVLA to new robotic setups, highlighting the effectiveness of pixel-level visual understanding and continuous action representation in PixelVLA. Notably, PixelVLA achieves significant performance in the LIBERO-Long setup, demonstrating its effectiveness in long-range manipulation.
@@ 5.3 New Robot Setups Adaptation Comparisons
Methods
Spatial Object Goal Long
Average
@@ 5.4 Ablation Studies
As presented in Tab. 3, incorporating the continuous action training stage (Baseline+FT+CAT) improves the average score of 3.8% compared to Baseline, highlighting the benefits of the proposed continuous action decoder. Further enhancement with pixel-level understanding (Baseline+FT+PUE) yields a more substantial gain of 8.0%. Ultimately, PixelVLA outperforms Baseline+FT+CAT by 6.3%. This progressive improvement validates the effectiveness of both pixel-level understanding and multimodal prompts in advancing visuomotor control capabilities.
@@ 6 Conclusion
This paper proposes PixelVLA, a vision-language-action (VLA) model, to address the limitations of existing VLAs, such as insufficient pixel-level understanding and over-reliance on textual prompts. PixelVLA integrates a multiscale pixel-aware encoder to inject pixel-level understanding, a continuous action decoder for generating accurate robotic actions, and a lightweight visual prompting encoder to support both textual and visual prompts. In addition, a two-stage automated annotation pipeline is designed to construct the Pixel-160K dataset containing 160K manipulation episodes. To advance fine-grained pixel-level understanding in VLAs, we propose a novel two-stage visuomotor instruction tuning framework to train PixelVLA, requiring only 1.5% of the pretraining cost of OpenVLA. Expensive evaluations on three VLA benchmarks show that PixelVLA can be integrated into existing VLAs to achieve a $10.1\%\sim 28.7\%$ improvement in manipulation success rate, effectively enhancing the spatial comprehension and complex environment adaptability of VLAs.

### v2

@@ 4.2 Visuomotor Tuning Data Generation
Gripper-aware Region Proposal Stage. Given a sequence of observations $\{\mathbf{x}^{1}_{\eta},\mathbf{x}^{2}_{\eta},\dots,\mathbf{x}^{N_{\eta}}_{\eta}\}$ from the $\eta$-th episode, the first gripper-close state in the episode as $G_{\eta}\in\{1,2,\dots,N_{\eta}\}$ and the corresponding observation as $\mathbf{x}^{G_{\eta}}_{\eta}\in\{\mathbf{x}^{1}_{\eta},\mathbf{x}^{2}_{\eta},\dots,\mathbf{x}^{N_{\eta}}_{\eta}\}$. Here, $N_{\eta}$ represents the length of the $\eta$-th episode. Sequentially, we can select a series of gripper-close state observations $\{\mathbf{x}^{G_{1}}_{1},\mathbf{x}^{G_{2}}_{2},\dots,\mathbf{x}^{G_{N_{e}}}_{N_{e}}\}$ from the whole dataset, where $N_{e}$ denotes the number of total episodes in the dataset. Furthermore, we assume $\{\mathbf{x}^{G_{1}}_{1},\mathbf{x}^{G_{2}}_{2},\dots,\mathbf{x}^{G_{N_{e}}}_{N_{e}}\}$ as a discrete video and apply SAM 2 (Ravi et al. (2024)) to generate $N_{e}$ gripper masks. Then, we compute the minimal axis-aligned bounding boxes enclosing these masks, uniformly enlarge each by a fixed margin to capture local context and reduce detection noise, and take the resulting boxes as the $N_{e}$ region proposals $\{\mathbf{R}_{1},\mathbf{R}_{2},\dots,\mathbf{R}_{N_{e}}\}$. Here, $\mathbf{R}_{\eta}\in\mathbb{R}^{4}$ is the proposal for the $\eta$-th episode. In this way, the region proposals can be leveraged to accurately capture object from cluttered and low-quality robot observations.
@@
Multimodal Object Segmentation Stage. Given a manipulation instruction such as “Put the Eggplant in Yellow Basket”, we employ Llama 2–7B to reason over the instruction and extract the textual description of the target object to be manipulated, e.g., “Eggplant”. For the $\eta$-th episode, we then provide the target object text along with the region proposal $\{\mathbf{R}_{\eta}\}$ into an open-vocabulary object detector Grounding DINO (Liu et al. (2024a)) and SAM (Kirillov et al. (2023)). These models detect all relevant object instances, generate their mask annotations, and associate them with the corresponding language expressions from the target object text. We then filter the predictions based on their confidence scores, retaining only the mask annotations within the bounding box that has the highest box-confidence. Sequentially, we derive visual prompts from the object masks by randomly sampling points within the mask, generating random lines inside the object region, and extracting external bounding boxes through mask contour detection.
Finally, we apply the proposed two-stage automated annotation pipeline to the publicly available Fractal dataset (Brohan et al. (2022)) and Bridge v2 dataset (Walke et al. (2023)). We first automatically discard samples with failed mask generation (e.g., empty or invalid masks) using a simple script, and then the authors rapidly inspect the remaining samples to remove those with clearly incorrect masks. In total, this process filters out approximately 19.2% of the generated samples. The resulting dataset, Pixel-160K, contains 160K robot manipulation episodes and 6.5M image–text–action triplets with visual prompts and mask annotations.
@@ 4.3 Visuomotor Instruction Tuning Procedure
To advance fine-grained pixel-level understanding in VLAs, we propose a novel visuomotor instruction tuning procedure, consisting of a continuous action training stage and a pixel-level understanding enhancement stage. Concretely, the first continuous action training stage enables the VLA model to acquire robust continuous action representations from a large mixture of image–text–action datasets. In the second stage, pixel-level understanding is explicitly enhanced by adapting the pretrained model on Pixel-160K dataset through LoRA adaptation (Hu et al. (2022); Li et al. (2025d)). The following sections elaborate on the key designs of this two-stage training strategy.
Table 1: SimplerEnv (Li et al. (2024c)) simulation evaluation results in terms of the average success rate for the Google Robot setup. VM denotes Visual Matching and VA is Variant Aggregation. ◼ and ◼ denote tuning-based methods applied to the pretrained weights of OpenVLA and $\pi_{0}$, respectively.
Methods Pick Coke Can Move Near Open/Close Drawer Average
@@
Continuous Action Training Stage. Before training, we initialize the vision encoder, the MLP projector, and the LLM backbone in PixelVLA with the pretrained weights of VLAs (Kim et al. (2024); Black et al. (2024)), which has been trained on the large-scale mixture dataset OXE (O’Neill et al. (2024)). In addition, during this stage, the visual prompt-aware encoder and the multiscale pixel-aware encoder of PixelVLA are removed, while all other modules except the continuous action decoder are frozen to preserve the general manipulation knowledge learned in the pretrained VLAs. To directly map the final hidden states of the last layer of LLM to continuous action values, we follow (Zhao et al. (2023); Kim et al. (2025)) to implement L1 regression to align predicted actions generated by the proposed continuous action decoder with the ground-truth actions. Unlike OpenVLA, which represents actions as discrete tokens by normalizing each action dimension to [1,+1][-1,+1] and uniformly discretizing it into 256 bins, PixelVLA directly predicts continuous action values, thereby avoiding the loss of fine-grained action details introduced by discretization. Furthermore, during this stage, we train PixelVLA on a mixture of Fractal dataset and Bridge v2 dataset.
Pixel-level Understanding Enhancement Stage. Originally, most existing visuomotor instruction tuning methods (Li et al. (2025b); Kim et al. (2025); Yang et al. (2025)) focus on image-level understanding. In contrast, at this stage, to enhance pixel-level understanding of PixelVLA, we employ LoRA adaptation to efficiently fine-tune PixelVLA’s LLM backbone on Pixel-160K dataset, while jointly training the visual prompt-aware encoder along with the multiscale pixel-aware encoder. Meanwhile, the continuous action decoder is optimized while the remaining PixelVLA modules remain frozen. Furthermore, we adopt the same L1 regression loss and continuous action representation strategy as those employed in the continuous action training stage. At each training step, given a mini-batch $\{\mathbf{x}^{i},\mathbf{p}^{i},\mathbf{a}^{i},\mathbf{L}^{i},\mathbf{V}^{i}\}^{B}_{i=1}$ sampled from the Pixel-160K dataset, the forward process at a single timestep of this stage can then be formulated as follows:
5 Experiments
@@ 5.1 Experimental Setup
Evaluation tasks. We conduct all experiments on three simulation benchmarks, i.e., SimplerEnv-Google Robot (Li et al. (2024c)), SimplerEnv-WidowX (Li et al. (2024c)) and LIBERO (Liu et al. (2023a)). SimplerEnv (Li et al. (2024c)) is an open-source simulation suite that facilitates reproducible and scalable evaluation of robot manipulation policies by explicitly addressing visual and dynamic gaps between simulation and real hardware. In light of this, we conduct zero-shot object manipulation comparisons on SimplerEnv. In addition, following OpenVLA (Kim et al. (2024)), we evaluate performance of new robot adaptation across four task suites within LIBERO (Liu et al. (2023a)), i.e., LIBERO-Spatial, LIBERO-Object, LIBERO-Goal and LIBERO-Long.
@@
Methods Put Spoon Put Carrot Stack Blocks Put Eggplant Average
@@
Implementation Details. To evaluate the effectiveness of PixelVLA, we apply its architecture and the proposed visuomotor instruction-tuning procedure to two widely-used VLAs, OpenVLA (Kim et al. (2024)) and $\pi_{0}$ (Black et al. (2024)). Regarding the training data, PixelVLA is trained in two stages: the first stage utilizes real-robot demonstrations from the Fractal dataset (Brohan et al. (2022)) and Bridge v2 dataset (Walke et al. (2023)), while the second stage employs 160K real-robot demonstrations from the proposed Pixel-160K dataset. For input robot observations across all datasets, PixelVLA is conditioned solely on a single third-person camera view and processes images at a resolution of 224×224 pixels. In all training stages, we set action chunk size to 8 for the continuous action decoder, i.e., the predicted action $\mathbf{a^{t}}\in\mathbb{R}^{8\times 7}$. The first training stage involves training PixelVLA for 100k steps with a batch size of 32 and a learning rate of $5\times 10^{-4}$. Notably, in light of the effectiveness action expert in $\pi_{0}$, we omit the first training stage when adapting PixelVLA on $\pi_{0}$. During the second training stage, we fine-tune the LLM backbone of PixelVLA using LoRA adaptation with a rank $r=32$. This stage is trained for 200k steps with a batch size of 32 and a learning rate of $1\times 10^{-3}$. In addition, to adapt PixelVLA to the LIBERO benchmark (Liu et al. (2023a)), we fine-tune the pre-trained model for 150K steps on each task suite using LoRA adaptation with rank $r=32$, a batch size of 32, and a learning rate of $5\times 10^{-4}$. In addition to the two baseline VLAs, OpenVLA (Kim et al. (2024)) and $\pi_{0}$ (Black et al. (2024)), we compare the performance of PixelVLA against other state-of-the-art VLAs, such as RT-1 (Brohan et al. (2022)), HPT (Wang et al. (2024)), Octo (Team et al. (2024)), TraceVLA (Zheng et al. (2024)), RoboVLMs (Liu et al. (2025)), Dita (Hou et al. (2025)) and SpatialVLA (Qu et al. (2025)). To further ensure fairness, we additionally fine-tune $\pi_{0}$ and OpenVLA on the Fractal and Bridge datasets, obtaining baselines denoted as $\pi_{0}$-SFT and OpenVLA-SFT.
@@
Methods Spatial Object Goal Long Average
@@
5.2 Zero-shot Object Manipulation Comparisons
This subsection evaluates the zero-shot manipulation performance of our model against baseline VLAs across multiple task categories and robot platforms. As shown in Tab. 1, on the Google Robot setup PixelVLA achieves an average VM score of 61.4 and VA score of 50.1, surpassing OpenVLA by 28.7/10.1 and OpenVLA-SFT by 29.8/11.5 in VM/VA, respectively. These results indicate a strong capability in both pixel-level understanding and adaptation to textual and visual prompts in out-of-domain adaptation. Notably, as shown in Fig. 4, PixelVLA outperforms TraceVLA and OpenVLA across various environmental variations, highlighting the effectiveness of the proposed visuomotor instruction tuning procedure in addressing out-of-domain generalization.
5.3 New Robot Setups Adaptation Comparisons
To evaluate the adaptability of PixelVLA to novel robotic setups and task configurations, we employ the proposed automated annotation pipeline to process the LIBERO benchmark training data (Liu et al. (2023a)), yielding the LIBERO-Pixel dataset. Subsequently, as summarized in Tab. 3, PixelVLA achieves state-of-the-art performance with an average success rate of 86.7 across all tasks, significantly surpassing strong baselines. In addition, As shown in Tab. 3, PixelVLA outperforms OpenVLA across all four LIBERO task suites, with particularly large gains on LIBERO-Long. We attribute this to the Continuous Action Decoder: action chunking helps the policy capture longer-range temporal dependencies (Liu et al. (2024b)), while continuous action prediction mitigates compounding discretization errors (Black et al. (2024)) over long-horizon manipulation. These superior results demonstrate the enhanced adaptability of PixelVLA to new robotic setups, highlighting the effectiveness of pixel-level visual understanding and continuous action representation in PixelVLA. Notably, PixelVLA achieves significant performance in the LIBERO-Long setup, demonstrating its effectiveness in long-range manipulation.
Table 4: Quantitative ablation studies on Variant Aggregation for the Google Robot setup, evaluated in the SimplerEnv simulation environment (Li et al. (2024c)).
Methods Pick Move Open/Close Average
 Coke Can  Near  Drawer
Baseline 54.5 47.7 17.7 40.0
       +FT 51.9 42.3 16.8 37.0
       +FT+CAT 61.3 52.3 17.7 43.8
       +FT+PUE 71.1 54.7 21.3 48.0
PixelVLA 72.7 57.7 20.0 50.1
@@ 5.4 Ablation Studies
As presented in Tab. 4, incorporating the continuous action training stage (Baseline+FT+CAT) improves the average score of 3.8% compared to Baseline, highlighting the benefits of the proposed continuous action decoder. Further enhancement with pixel-level understanding (Baseline+FT+PUE) yields a more substantial gain of 8.0%. Compared to the single-stage Baseline+FT+PUE, PixelVLA adds a continuous action training stage, and this two-stage scheme leads to a slight drop on Open/Close Drawer, due to the catastrophic forgetting (Li et al. (2024a)) in joint two-stage optimization and the high difficulty and sensitivity of this task. Ultimately, PixelVLA outperforms Baseline+FT+CAT by 6.3%. This progressive improvement validates the effectiveness of both pixel-level understanding and multimodal prompts in advancing generalization capabilities.
@@ 6 Conclusion
This paper proposes PixelVLA, a vision-language-action (VLA) model, to address the limitations of existing VLAs, such as insufficient pixel-level understanding and over-reliance on textual prompts. PixelVLA integrates a multiscale pixel-aware encoder to inject pixel-level understanding, a continuous action decoder for generating accurate robotic actions, and a lightweight visual prompt-aware encoder to support both textual and visual prompts. In addition, a two-stage automated annotation pipeline is designed to construct the Pixel-160K dataset containing 160K manipulation episodes. To advance fine-grained pixel-level understanding in VLAs, we propose a novel two-stage visuomotor instruction tuning framework to train PixelVLA, requiring only 1.5% of the pretraining cost of OpenVLA. Expensive evaluations on three VLA benchmarks show that PixelVLA can be integrated into existing VLAs to achieve a $10.1\%\sim 28.7\%$ improvement in manipulation success rate, effectively enhancing the spatial comprehension and complex environment adaptability of VLAs.
Acknowledgments
This work is supported by the National Key Research and Development Program of China (2023YFB4704800) and National Nature Science Foundation of China under Grant (62225310,62273333), and the Fundamental Research Funds for the Central Universities (2025ZYGXZR032, 2024ZYGXZR024).